OpenStudy (anonymous):

The equation y = 16x + 560 describes the number of calories in a recipe, where y is the total number of calories and x is the number of teaspoons of sugar added. The table shows the number of calories in another recipe, where y is the total number of calories and x is the number of teaspoons of honey added. How many calories are in a teaspoon of sugar, and how many are in a teaspoon of honey? Teaspoons of honey 1 2 4 7 Total calories 504 568 696 888 A teaspoon of sugar has____calories. A teaspoon of honey has____calories.

OpenStudy (anonymous):

Sugar 16 and Honey 64 ?

OpenStudy (jmark):

let the equation for honey be y = ax + b where a = spoons of honey. Putting values from table we get, 504 = a + b and 568 = 2a + b. Solving we have a = 64 and b = 440, equation is y = 64x + 440. Then sugar has 16 calories in a teaspoon and a teaspoon honey has 64 calories.